Add and Simplify
2x/x+3+5/x-3=

Answer by fractalier(2101)

If you mean
2x/(x+3) + 5/(x-3) =
then we change each fraction to its lowest common denominator, or (x+3)(x-3) and get
2x(x-3) / (x+3)(x-3) + 5(x+3) / (x+3)(x-3) =
2x^2 - 6x + 5x + 15 / (x+3)(x-3) =
2x^2 - x + 15 / (x+3)(x-3)